Let X be a random variable . P(X=-2)=P(X=-1),P(X=2)=P(X=1),P(X=0)=P(X>0)=P(X<0) . obtain the probability mass function of X


1
Expert's answer
2021-05-28T10:00:54-0400

Probability mass function:


"P(X = 0) = P(X>0) = P(X<0) = \dfrac{1}{3}"


"P(X=-2) = P(X = -1) = \dfrac{1}{6}"


"P(X=2) = P(X=1) = \dfrac{1}{6}"